OpenStudy (anonymous):

Choose the procedure you would use to solve the equation x/3 − 4 = 13 for x. Add four to both sides, and then divide both sides by three. Add four to both sides, and then multiply both sides by three. Multiply both sides by three, and then subtract both sides by four. Divide both sides by three, and then add four to both sides.

OpenStudy (anonymous):

Add four, now you would get x/3 = 13 + 4

OpenStudy (dumbsearch2):

Choose the procedure you would use to solve the equation x over 3 − 4 = 13 for x. ---B)Add four to both sides, and then multiply both sides by three.

OpenStudy (dumbsearch2):

Thats the answer, I'm pretty sure. You can verify that, of course. ^_^

OpenStudy (anonymous):

you dont add 4 to both sides